OpenStudy (meilendurcer):

Which ratio forms a proportion with 9/15 ? A. 3/6 B. 2/3 C. 12/30 D. 6/10

OpenStudy (anonymous):

9/15=3/5 6/10=3/5....hence they are in proportion

OpenStudy (phi):

notice that you can simplify 9/15 by dividing top and bottom by 3

OpenStudy (meilendurcer):

Which proportion has cross products of 5 × 24 and 8 × 15? A. 5/8=15/24 B. 5/24= 8/15 C. 8/24/= 15/5 D. 15/8= 5/24

OpenStudy (phi):

cross product means multiply the top of one side times the bottom of the other side

OpenStudy (meilendurcer):

is it C?

OpenStudy (phi):

in C, 8 is on the top on the left. you multiply it times the bottom on the right side. 5 is on the bottom on the right. that means 8*5 is that what you want?

OpenStudy (meilendurcer):

im not sure

OpenStudy (meilendurcer):

@pkjha3105 do you know what it is?

OpenStudy (phi):

look at each choice: write down the top of left side times the bottom of the right side can you do that? then write down the bottom of the left side times the top of the right side. can you do that?

OpenStudy (meilendurcer):

75 and 192

OpenStudy (phi):

don't multiply them out, just write them down example A. 5/8=15/24 top left times bottom right --> 5*24 bottom left times top right --> 8*15 Is that what you want? Which proportion has cross products of 5 × 24 and 8 × 15?

OpenStudy (meilendurcer):

yes.

OpenStudy (phi):

Here is choice B B. 5/24= 8/15 top left times bottom right --> 5*15 bottom left times top right ---> 24*8 is that what you want? NO

OpenStudy (meilendurcer):

i am not getting this.

OpenStudy (phi):

You have to answer the question Which proportion has cross products of 5 × 24 and 8 × 15? A. 5/8=15/24 choice A gives you